# K-theory and localizing invariants of large categories

### v2

@@ 1.17. Products of dualizable categories
In this subsection we discuss the products in It turns out that they are very well-behaved: the functor commutes with products, and the category satisfies the weak (AB4*) axiom and the (AB6) axiom.
First, we make an observation about infinite products of ind-objects. Let be a family of small stable categories. Given ind-objects we can consider each as an object of We denote by the product of these objects in We denote by the resulting functor. This is simply the right adjoint to the natural functor We give two more equivalent descriptions of the functor
@@
Let be a family of homological epimorphisms between small stable categories.
1) The functor is a homological epimorphism.
2) For a family of objects the object is contained in if and only if each is contained in Moreover, the category is generated by the objects of this form.
@@
Put
1) Take some families of objects Since are homological epimorphisms, using Proposition 1.101, we obtain equivalences
Hence, is a homological epimorphism.
2) The first statement holds even without the assumption that are homological epimorphisms, it follows from Corollary 1.100. For the “moreover” statement, note that the category is generated by the objects of the form
@@
Let be a collection of dualizable categories.
1) We have a short exact sequence
@@
By Theorem 1.92, we have an equivalence
Now part 1) follows from Proposition 1.102, applied to the homological epimorphisms
Part 2) follows from the fact that an infinite product of short exact sequences in is a short exact sequence, applied to the sequences
@@
The image of contains all objects of the form where and is the evaluation functor for Moreover, the image of is generated by the objects of this form.
2) Let be a family of strongly continuous functors between dualizable categories. Then the following square commutes:
Here the functors and are as in 1). The functor (resp. ) is (resp. ).
@@
1) The functor is obtained by applying to the inclusion in the short exact sequence (1.9). Note that the (fully faithful) functor takes to the functor The assertion now follows from Proposition 1.102, part 2), applied to the homologicl epimorphisms
2) We need to show that the image of is contained in the image of By part 1), it suffices to show that for any the object is contained in the image of We have
@@
The only statement that requires a proof is that the functor is a quotient functor. Equivalently, it suffices to show that the functor is a quotient functor, i.e. its right adjoint is fully faithful. By Lemma 1.104, the latter functor is the upper horizontal arrow in a commutative square
@@
We first observe that (AB6) holds in the category of small stable Karoubi-complete categories, since the category is compactly generated (one can also check (AB6) for directly, by showing that the comparison functor is fully faithful and essentially surjective). Since the functor commutes both with products and with filtered colimits, we deduce that (1.10) is an equivalence when all are compactly generated.
By Propositions 1.68 and 1.105, filtered colimits and products of short exact sequences in are short exact. Consider the map of short exact sequences
@@ 2. Dualizability is equivalent to flatness
In this section we give a non-trivial criterion of dualizability of a presentable stable category via a “flatness” condition. It can be skipped on the first reading.
Let be a presentable stable category. We can ask the following naive question: is it true that for any fully faithful colimit-preserving functor between presentable stable categories, the functor is also fully faithful?
@@
A category is called flat if the answer to the above question is affirmative, i.e. the tensor product preserves fully faithful functors (=monomorphisms).
We have the following surprising result.
@@
Let be a presentable stable category. Then is flat if and only if it is dualizable.
The “if” direction is easy: dualizability of implies an equivalence To prove the “only if” direction, we need a certain “extra -functoriality” statement.
Denote by the -category of presentable stable categories and exact accessible functors. We have the following result.
@@
Let be a presentable stable category. Then we have a natural oplax -functor which induces the usual Lurie tensor product when we restrict to continuous functors. Moreover, if is flat, then we get an actual -functor
@@
Givem presentable stable categories and let us denote by the -category of triples where and are fully faithful continuous functors, such that we have a semi-orthogonal decomposition
@@
Then is in fact an -category, and it is naturally equivalent to the category of exact accessible functors Namely, given a triple the corresponding functor is given by where is the right adjoint to Conversely, to an exact accessible functor we associate the category as in Definition 1.1. Recall that is the category of triples where and The inclusions and are given respectively by and
Given correspondences and we can define their composition as follows. We form the pushout (in ), and take the full subcategory generated by the images of the compositions and Then is naturally an object of Moreover, this composition operation is compatible with the composition of the corresponding accessible functors.
Now, we observe that for any presentable and we have a well-defined functor Indeed, given we have Here the fully faithfulness of (resp. ) follows from the existence of a continuous right adjoint to (resp. a left adjoint to ).
Now, given and as above, we observe that we have a natural equivalence Thus, we have a natural transformation
in Moreover, if is flat, then (2.1) is an equivalence.
Identifying the correspondences with accessible functors, we obtain an oplax -functor If is flat, we get an actual -functor.
Finally, if the triple corresponds to a continuous functor then the functor is continuous. Hence, the functor is the right adjoint to and the object corresponds to the functor – the usual Lurie tensor product.∎
Proof of Theorem 2.2..
We already observed that dualizability implies flatness by identifying with
Now, suppose that is flat. Using Proposition 2.3 (extra -functoriality), we will deduce that (AB6) holds in By Proposition 1.54 this will imply that is dualizable.
Consider a family of filtered categories We define the following presentable (in fact, compactly generated) stable categories:
We have the following natural exact accessible functors:
- the product of the colimit functors
- the product functor;
- taking to the functor
- the colimit functor.
Since the axiom (AB6) holds in the category the natural transformation is an equivalence. So we have a commutative square
Applying the extra 2-functoriality of we obtain a commutative square
But it is straightforward to check that the latter square is of the form
Moreover, the functors in this commutative square are analogous to the functors with “” replaced by “”. Therefore, commutativity of (2.2) exactly means that (AB6) holds in ∎
@@ 3. Extensions of compactly generated categories
It turns out that in general the answer is “no”. To give a precise criterion of dualizability of we introduce some notation.
Denote by the right adjoint to and by the right adjoint to Note that we have a semi-orthogonal decomposition Putting we obtain an equivalence
Note that we have equivalences of categories of functors:
We denote by the functor corresponding to
Recall the following full subcategories of formed by (elementary) Tate objects.
@@
Assuming Proposition 3.3, we deduce the following result.
@@
The special case of Theorem 3.4 for split extensions is well-known. Namely, short exact sequences of the form (3.2) where both functors have right (resp. left) adjoints correspond to semi-orthogonal decompositions of the form (resp. ), which in turn correspond to functors (resp. ), i.e. -bimodules (resp. -bimodules).
@@
Clearly, .
If is dualizable, then for any object there exists an object such that . Such objects together with the objects of form a generating collection of compact objects of hence is compactly generated.
. We use the above observation: is compactly generated if and only if for any object there exists a compact object such that Via the identification the object must be of the form
We need to describe the necessary and sufficient conditions for to be a compact object of Note that commutes with colimits since is a compact object of We have
We deduce that the object is compact if and only if the right-hand side in (3.3) commutes with coproducts.
Note that the functor is exactly the object where corresponds to as above. We conclude that the existence of such that is equivalent to the existence of objects and and an exact triangle
@@
We give some natural examples of non-dualizable categories which are extensions of compactly generated categories. We use the notation for the direct sum of copies of an object indexed by the elements of a set
@@
Consider the functor corresponding to By Proposition 3.3, we need to show that
Identifying with the opposite of the category of -linear accessible functors we see that corresponds to the functor Using the fact that (AB6) holds in we obtain isomorphisms
Putting we obtain an isomorphism
To show that we need to show that the image of in is not a pro-constant object. Assume the contrary. Since the maps are split epimorphisms for we deduce that the images must form an eventually constant pro-system in But this is clearly false: for any the map is not an isomorphism in since ∎
@@
Then Theorem 3.10 implies that is not dualizable. Indeed, the corresponding subset contains the ideals and but does not contain the ideal
@@
Denote by the injective hull of the -module Then is a locally -torsion -module, hence Moreover, is an injective cogenerator of the abelian category of -modules.
We first observe that has bounded cohomology. Indeed, suppose that for infinitely many integers For each choose a non-zero map (which exists since is a cogenerator). Since is injective, we obtain the map This map does not factor through a finite direct sum, which contradicts the commutation of with coproducts.
Next, we show that each -module is finitely generated. We may and will assume that Suppose that is not finitely generated. Then we can construct a strictly increasing sequence of submodules
such that each is finitely generated. Put For choose a non-zero map such that (where ). Choose some which extends for each The sequence gives a well-defined map which does not factor through a finite direct sum. Since is noetherian, the module is injective. Choosing an extension of we obtain a morphism which does not factor through a finite direct sum. This contradicts the commutation of with coproducts.
So we showed that has bounded finitely generated cohomology. To prove that is perfect, it suffices to observe that is perfect over by adjunction, the functor commutes with coproducts. ∎
@@
where and are strongly continuous, and the categories on the left and on the right are compactly generated.
Suppose that is a dualizable category. The proof of Proposition 3.3 shows that there exists an object and a map such that the functor
@@
Suppose that is convex. Denote by the set of all points which are specializations of elements of Put Then and are closed under the specialization, hence the categories and are compactly generated and the inclusion is strongly continuous. Hence, the quotient is also compactly generated.
Now let be a non-convex subset. Suppose that the category is dualizable. Choose some such that and Since is noetherian, we may and will assume that is a minimal element of the set Denote by and respectively the prime ideals corresponding to the points and Put We identify the set with its image in and put
Since the category is dualizable, so is the category Moreover, since is dualizable, the functor
@@
Then where is the maximal ideal, and the field of fractions of is
Consider the non-convex subset Then we have
– a compactly generated category.
To see this, it suffices to observe that if is an -torsion -module, then acts by zero on Indeed, if is an element such that then It follows that for we have
@@ 4.1. Universal localizing invariants
To avoid set-theoretic issues, we will consider only accessible localizing invariants with values in accessible stable categories.
Let be an accessible stable category with (not necessarily cocomplete). Recall that a functor is a localizing invariant if the following conditions hold:
@@
If is a regular cardinal such that has -filtered colimits, we denote by the category of localizing invariants which commute with -filtered colimits. It follows from [BGT13] that for any regular cardinal there exists a universal localizing invariant such that the category is accessible and has -filtered colimits, and for any other such category we have
Here the right-hand side is the category of exact functors which commute with -filtered colimits.
A clarification is required when is uncountable. In this case the construction of [BGT13, Section 8.2] gives a solution to a different universal problem: the authors construct a presentable stable category and a localizing invariant which commutes with -filtered colimits, such that for any presentable stable category we have
@@
Any short exact sequence in is an -filtered colimit of short exact sequenes of -compact categories.
@@
in Choose an -directed system such that each is in and Consider a poset of pairs where and is a full stable idempotent-complete subcategory generated by at most countable collection of objects. By 1.6, such a category is automatically -compact. The poset is -directed, and we conclude that (4.1) is a -indexed colimit of short exact sequences
@@
Recall that for each regular the category has a natural (closed) symmetric monoidal structure, and the functor is symmetric monoidal. Below we will often formulate results in terms of the universal localizing invariants. They automatically imply the corresponding results for general localizing invariants.
@@
In the forthcoming papers [E1, E2] we will study the categories and in great detail. In particular, we will prove that the category and its relative versions are rigid as a symmetric monoidal categories in the sense of Gaitsgory and Rozenblyum [GaRo17, Definition 9.1.2]. This in particular will imply that the category is dualizable, which is quite non-obvious from the definition. In this paper however we will use only the universal property of the categories

### v3

@@ 1.17. Products of dualizable categories
In this subsection we discuss the products in It turns out that they are very well-behaved: the functor commutes with products, and the category satisfies the weak (AB4*) axiom and the (AB6) axiom. First, we make an observation about infinite products of ind-objects. Let be a family of small stable categories. Given ind-objects we can consider each as an object of We denote by the product of these objects in We denote by the resulting functor. This is simply the right adjoint to the natural functor We give another two equivalent descriptions of the functor
@@
Let be a family of homological epimorphisms between small stable categories. 1) The functor is a homological epimorphism. 2) For a family of objects the object is contained in if and only if each is contained in Moreover, the category is generated by the objects of this form.
@@
Put 1) Take some families of objects Since are homological epimorphisms, using Proposition 1.101, we obtain equivalences
Hence, is a homological epimorphism. 2) The first statement holds even without the assumption that are homological epimorphisms, it follows from Corollary 1.100. For the “moreover” statement, note that the category is generated by the objects of the form
@@
Let be a collection of dualizable categories. 1) We have a short exact sequence
@@
The image of contains all objects of the form where and is the evaluation functor for Moreover, the image of is generated by the objects of this form. 2) Let be a family of strongly continuous functors between dualizable categories. Then the following square commutes:
Here the functors and are as in 1). The functor (resp. ) is (resp. the right adjoint to ).
@@
1) The functor is obtained by applying to the inclusion in the short exact sequence (1.9). Note that the (fully faithful) functor takes to the functor The assertion now follows from Proposition 1.102, part 2), applied to the homological epimorphisms 2) We need to show that the image of is contained in the image of By part 1), it suffices to show that for any the object is contained in the image of We have
@@
The only statement that requires a proof is that the functor is a quotient functor. Equivalently, we need to show that the dual functor is fully faithful. By Lemma 1.104, the latter functor is the upper horizontal arrow in a commutative square
@@
We first observe that (AB6) holds in the category of small stable Karoubi-complete categories, since the category is compactly generated (one can also check (AB6) for directly, by showing that the comparison functor is fully faithful and essentially surjective). Since the functor commutes both with products and with filtered colimits, we deduce that (1.10) is an equivalence when all are compactly generated. By Propositions 1.67 and 1.105, filtered colimits and products of short exact sequences in are short exact. Consider the map of short exact sequences
@@ 2. Dualizability is equivalent to flatness
In this section we give a non-trivial criterion of dualizability of a presentable stable category via a “flatness” condition. It can be skipped on the first reading. Let be a presentable stable category. We can ask the following naive question: is it true that for any fully faithful colimit-preserving functor between presentable stable categories, the functor is also fully faithful?
@@
1) A category is called flat in if the answer to the above question is affirmative, i.e. the functor preserves fully faithful functors (=monomorphisms). 2) Let be a regular cardinal. A -presentable stable category is called flat in if the functor preserves fully faithful functors.
Note that the flatness in is a non-trivial condition only if is uncountable. We have the following surprising result.
@@
Let be an uncountable regular cardinal. Then a -presentable stable category is flat in if and only if satisfies (AB6) for -small products. In particular, a presentable stable category is flat in if and only if it is dualizable.
The “if” direction in the second statement is easy: dualizability of implies an equivalence The “if” direction in the first statement is non-trivial, see below for the proof. Our argument for the “only if” direction will use a certain “extra -functoriality” statement. After the previous version of the paper was published, German Stefanich found a different argument, which in fact works in a much more general relative context (and for not necessarily stable categories). We will use the model for -categories given by complete Segal space objects of the -category of (possibly very large) -categories. We refer to [Lur09b, Section 1] for a detailed discussion. More precisely, we consider -categories as cartesian fibrations of -categories satisfying the following conditions.
Denote by the fiber of over Then for the natural functor
(the -fold fiber product) is an equivalence of -categories (the Segal condition).
The -category is a discrete groupoid.
Given -categories an oplax -functor is a functor between fibered categories over If moreover is cartesian (i.e. preserves cartesian maps), then is an actual -functor. From now on, we fix an uncountable regular cardinal Denote by the -category of -presentable stable categories and exact -accessible functors. It is convenient to identify with the following (very large) cartesian fibration over Given -presentable stable categories we denote by the category of cartesian fibrations together with equivalences such that the transition functors are exact -accessible functors. Then the fiber over is given by the disjoint union of categories over all -tuples Given objects in and in the fiber of the mapping space over an order-preserving map is non-empty only if for and in this case it is the space of functors over which are compatible with the equivalences We denote by the locally full sub -category which is given by the full subcategories formed by cartesian fibrations with continuous transition functors. We have the following result.
@@
Let be a -presentable stable category. Then we have a natural oplax -functor which induces the usual Lurie tensor product Moreover, if is flat in then we have an actual -functor
To prove this result, we need the following statement about semi-orthogonal decompositions in
Lemma 2.4.
Let be -presentable stable categories, and suppose that we have fully faithful -strongly continuous functors which give a semi-orthogonal decomposition
Then the functors are also fully faithful, and they give a semi-orthogonal decomposition
@@
It suffices to prove this for Note that the functor has a left adjoint and the functor has a continuous right adjoint Hence, the functors and are fully faithful. Their images generate by colimits. Finally, to see the the semi-orthogonality, note that is the left adjoint to and we have
This is equivalent to the required semi-orthogonality. ∎
Proof of Proposition 2.3..
@@
Given -presentable stable categories and let us denote by the -category of triples where is a -presentable stable category and are fully faithful -strongly continuous functors, such that we have a semi-orthogonal decomposition
@@
Then the category is naturally equivalent to the category of exact -accessible functors Namely, given a triple the corresponding functor is given by where is the right adjoint to Conversely, to an exact accessible functor we associate the category as in Definition 1.1. Recall that is the category of triples where and The inclusions and are given respectively by and We now formally define the -category as a (very large) cartesian fibration over satisfying the Segal condition, such that the fiber over is the (very large) set of -presentable stable categories. Given -presentable stable categories where we define the -category to be the category of -tuples where is a -presentable stable category and are fully faithful -strongly continuous functors such that the following conditions hold:
the functors give a semi-orthogonal decomposition
if we denote by the right adjoint to then we have isomorphisms
If we define to be a single point, considered as an -category. We think of this point as a pair The condition (2.1) (which is non-trivial only for ) can be reformulated as follows (assuming that we have the required semi-orthogonal decomposition). Denote by the full stable subcategory generated by and Then (2.1) holds if and only if we have an equivalence
The fiber over is given by the disjoint union of categories over all -tuples The fiber of the mapping space
over an order-preserving map map is the following. It is non-empty only if for and in this case it is the space of functors together with isomorphisms It is easy to see that is a cartesian fibration, and it follows from (2.2) that this fibration satisfies the Segal condition. Now, we claim that we have a natural equivalence of -categories which induces the identity map on objects. The functor sends to the following cartesian fibration The fiber over is given by and for the mapping space is given by
The inverse functor sends a cartesian fibration to the category of sections of the dual cocartesian fibration (with the same fibers and the same transition functors). Here for the functor is the composition of with the left adjoint to the evaluation at It is easy to see that the functors can be naturally extended to a functor between the cartesian fibrations over which preserves cartesian maps. Thus, we get an equivalence of -categories, as stated. We claim that for a -presentable stable category we have a natural functor between the cartesian fibrations over which does not necessarily preserve cartesian maps. On it is given by Given a sequence we have a functor
given by
Note that has the required semi-orthogonal decomposition by Lemma 2.4. Also, by loc. cit. the functors are fully faithful. Since satisfies the condition (2.2) and since the functor commutes with pushouts, we see that is indeed an object of We obtain an oplax -functor Using the above equivalence we obtain an oplax -functor We claim that it induces the usual Lurie tensor product Indeed, the cartesian fibration (corresponding to ) is the full subfibration of which consists of such that each functor is strongly continuous. Thus, the oplax -functor induces an actual -functor which corresponds to the usual Lurie tensor product. Finally, we need to show that if is flat in then we get an actual -functor We need an explicit description of the composition of -morphisms in Given correspondences and their composition can be obtained as follows. We form the pushout
and take the full subcategory generated by the images of the compositions and Then is naturally an object of and it is the composition Now, the flatness of implies that the functor
is fully faithful. Hence, the map
in is an isomorphism. Therefore, is an actual -functor, as required. ∎
Note that the oplax -functor can be described on -morphisms more explicitly. For a -presentable stable category we identify with the category of -left exact functors (these are functors which commutes with -small limits). Then for an exact -accessible functor the functor is identified with the composition
where is the left adjoint to the inclusion. However, with this description it is not clear how to show that flatness of in implies that we get an actual -functor.
Proof of Theorem 2.2.
By Proposition 1.53, the second assertion follows from the first one. As above we fix an uncountable regular cardinal and consider a -presentable stable category For brevity we say that satisfies (AB6)κ if (AB6) for -small products holds in First suppose that satisfies (AB6) Take some fully faithful -strongly continuous functor between -presentable stable categories. The right adjoint to is identified with the precomposition functor
We need to show that this is a quotient functor. By the proof of Proposition 1.53, our assumption on means that the colimit functor commutes with -small limits. Its right adjoint commutes with all limits, hence the functor (2.3) is a retract of the functor
The latter is a quotient functor since the category is dualizable, hence flat. The class of accessible quotient functors is closed under retracts, hence (2.3) is a quotient functor, as required. Now, suppose that is flat in Using Proposition 2.3 (extra -functoriality), we will deduce that (AB6)κ holds in Consider a -small family of filtered categories We have the following natural exact -accessible functors between -presentable stable categories: - the product of the colimit functors - the product functor; - taking to the functor - the colimit functor. The functors and are continuous, and the functors and have left adjoints and respectively. Here is the diagonal functor , and the components of are the left Kan extension functors Tautologically, we have a commutative square
Now, (AB6)κ for means that the square (2.4) satisfies the dual Beck-Chevalley condition: the induced map is an isomorphism. Note that the square (2.4) is equivalent to the following tensor product:
Since the axiom (AB6) holds in the category the map is an isomorphism. Applying the -functoriality of given by Proposition 2.3 (since is flat in ), we conclude that the map is an isomorphism. Therefore, (AB6)κ holds in as required. ∎
@@ 3. Extensions of compactly generated categories
It turns out that in general the answer is “no”. To give a precise criterion of dualizability of we introduce some notation. Denote by the right adjoint to and by the right adjoint to Note that we have a semi-orthogonal decomposition Putting we obtain an equivalence Note that we have equivalences of categories of functors:
We denote by the functor corresponding to Recall the following full subcategories of formed by (elementary) Tate objects.
@@
Assuming Proposition 3.3, we deduce the following result, which first appeared in [CDH+20, Appendix A].
@@
The special case of Theorem 3.4 for split extensions is well known. Namely, short exact sequences of the form (3.2) where both functors have right (resp. left) adjoints correspond to semi-orthogonal decompositions of the form (resp. ), which in turn correspond to functors (resp. ), i.e. -bimodules (resp. -bimodules).
@@
Clearly, . . If is dualizable, then by Proposition 1.18 for any object there exists an object such that . Such objects together with the objects of form a generating collection of compact objects of hence is compactly generated. . We use the above observation: is compactly generated if and only if for any object there exists a compact object such that Via the identification the object must be of the form We need to describe the necessary and sufficient conditions for to be a compact object of Note that commutes with colimits since is a compact object of We have
We deduce that the object is compact if and only if the right-hand side in (3.3) commutes with coproducts. Note that the functor is exactly the object where corresponds to as above. We conclude that the existence of such that is equivalent to the existence of objects and and an exact triangle
@@
We give a natural “linear-algebraic” example of a non-dualizable category which is an extension of compactly generated categories. We use the notation for the direct sum of copies of an object indexed by the elements of a set
@@
Consider the functor corresponding to By Proposition 3.3, we need to show that Identifying with the opposite of the category of -linear accessible functors we see that corresponds to the functor Using the fact that (AB6) holds in we obtain isomorphisms
Putting we obtain an isomorphism To show that we need to show that the image of in is not a pro-constant object. Assume the contrary. Since the maps are split epimorphisms for we deduce that the images must form an eventually constant pro-system in But this is clearly false: for any the map is not an isomorphism in since ∎
@@
Then Theorem 3.10 implies that is not dualizable. Indeed, the corresponding subset of is given by We see that contains the ideals and but does not contain the ideal hence is not convex.
@@
Denote by the injective hull of the -module Then is a locally -torsion -module, hence Moreover, is an injective cogenerator of the abelian category of -modules. We first observe that has bounded cohomology. Indeed, suppose that for infinitely many integers For each choose a non-zero map (which exists since is a cogenerator). Since is injective, we obtain the map This map does not factor through a finite direct sum, which contradicts the commutation of with coproducts. Next, we show that each -module is finitely generated. We may and will assume that Suppose that is not finitely generated. Then we can construct a strictly increasing sequence of submodules
such that each is finitely generated. Put For choose a non-zero map such that (where ). Choose some which extends for each The sequence gives a well-defined map which does not factor through a finite direct sum. Since is noetherian, the module is injective. Choosing an extension of we obtain a morphism which does not factor through a finite direct sum. This contradicts the commutation of with coproducts. So we showed that has bounded finitely generated cohomology. To prove that is perfect, it suffices to observe that is perfect over by adjunction, the functor commutes with coproducts. ∎
@@
where and are strongly continuous, and the categories on the left and on the right are compactly generated. Suppose that is a dualizable category. The proof of Proposition 3.3 shows that there exists an object and a map such that the functor
@@
Suppose that is convex. Denote by the set of all points which are specializations of elements of Put Then and are closed under the specialization, hence the categories and are compactly generated and the inclusion is strongly continuous. Hence, the quotient is also compactly generated. Now let be a non-convex subset. Suppose that the category is dualizable. Choose some such that and Since is noetherian, we may and will assume that is a minimal element of the set Denote by and respectively the prime ideals corresponding to the points and Put We identify the set with its image in and put Since the category is dualizable, so is the category Moreover, since is dualizable, the functor
@@
Then where is the maximal ideal, and the field of fractions of is Consider the non-convex subset Then we have
– a compactly generated category. To see this, it suffices to observe that if is an -torsion -module, then acts by zero on Indeed, if is an element such that then It follows that for we have
@@ 4.1. Universal localizing invariants
To avoid set-theoretic issues, we will consider only accessible localizing invariants with values in accessible stable categories. Let be an accessible stable category (not necessarily cocomplete). Recall that a functor is a localizing invariant if the following conditions hold:
@@
If is a regular cardinal such that has -filtered colimits, we denote by the category of localizing invariants which commute with -filtered colimits. It follows from [BGT13] that for any regular cardinal there exists a universal localizing invariant such that the category is accessible and has -filtered colimits, the functor is -continuous and for any other such category we have
Here the right-hand side is the category of exact functors which commute with -filtered colimits. A clarification is required when is uncountable. In this case the construction of [BGT13, Section 8.2] gives a solution to a different universal problem: the authors construct a presentable stable category and a localizing invariant which commutes with -filtered colimits, such that for any presentable stable category we have
@@
Let be an uncountable regular cardinal. Any short exact sequence in is a -filtered colimit of short exact sequences of -compact categories.
@@
in Choose a -directed system such that each is in and Consider a poset of pairs where and is a full stable idempotent-complete subcategory generated by a -small collection of objects. By Proposition 1.6, such a category is automatically -compact. The poset is -directed, and we conclude that (4.1) is a -indexed colimit of short exact sequences
@@
Recall that for each regular cardinal the category has a natural symmetric monoidal structure, and the functor is symmetric monoidal. Below we will often formulate results in terms of the universal localizing invariants. They automatically imply the corresponding results for general localizing invariants.
@@
In the forthcoming papers [E1, E2] we will study the categories and in great detail. In particular, we will prove that the category and its relative versions are rigid as a symmetric monoidal categories in the sense of Gaitsgory and Rozenblyum [GaiRoz17, Definition 9.1.2]. This in particular will imply that the category is dualizable, which is quite non-obvious from the definition. In this paper however we will use only the universal property of the categories